Ask
your own question, for FREE!
Mathematics
12 Online
Audrey has a bag of similar-sized red and green grapes. There are 12 red grapes and 13 green grapes in the bag. If the first 7 grapes she ate were green what is the probability that the 8th grapes she eats is green and the 9th grape she eats is red?
Still Need Help?
Join the QuestionCove community and study together with friends!
P(g) * P(r) Probability of it being green at 8th is (13-7)/(12+13 - 7) = 6/18 = 1/3 Probability of it being red at 9th is 12/(12+13 - 8) = 12/17 P(g) * P(r) = 1/3 * 12/17 = 12/51 = 4/17
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
5 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal